Circular No. 6125

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

COMET C/1995 A1 (JEDICKE)

The following precise positions have been reported:

1995 UT α_{2000} δ_{2000} m_{1} Observer
Jan. 9.19482 7^{h}55^{m}45.^{s}75 +20^{o}24′13′′.8 19.1 Jedicke
9.20357 7 55 45.44 +20 24 15.9  _{”}
9.21167 7 55 45.08 +20 24 17.9  _{”}
10.56418 7 54 58.44 +20 30 06.6 18.3 Kojima
10.59659 7 54 57.39 +20 30 15.3  _{”}

R. Jedicke (Kitt Peak). 0.9-m Spacewatch telescope. T. Kojima (Chiyoda). 0.25-m f/6 reflector. Communicated by S. Nakano.

The parabolic orbital elements below are extremely indeterminate, but the ephemeris should aid further observations. It is rather probable that the comet is of short period.

T = 1994 Apr. 13.91 TT  ω = 294.^{o}14
Ω = 116.07 }2000.0
q = 2.7624 AU  i = 12.38
1995TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Jan. 3 7^{h}59.^{m}24 +19^{o}57.^{′}6 2.869 3.825 164.^{o}3 4.^{o}0 19.1
8 7 56.44 +20 19.1 2.885 3.858 170.1 2.5 19.2
13 7 53.57 +20 40.5 2.909 3.891 176.0 1.0 19.2
18 7 50.70 +21 01.4 2.941 3.924 178.2 0.4 19.3
23 7 47.90 +21 21.7 2.980 3.957 172.4 1.9 19.3
28 7 45.24 +21 40.9 3.026 3.991 166.7 3.3 19.4
Feb. 2 7 42.80 +21 59.0 3.081 4.025 161.0 4.6 19.5
7 7 40.63 +22 15.8 3.142 4.059 155.3 5.8 19.6
12 7 38.77 +22 31.1 3.210 4.093 149.8 7.0 19.7

EF PEGASI P. Schmeer, Bischmisheim, Germany, reports his detection of a rare

outburst of this SU UMa-type dwarf nova (cf. IAUC 5369, 5371, 5372, 5378). Visual magnitude estimates: Jan. 5.73 UT, [13.0 (Schmeer); 10.99, [13.5 (J. E. Bortle, Stormville, NY); 11.73, 11.1 (Schmeer); 11.74, 11.1 (G. Poyner, Birmingham, England); 12.70, 10.8 (L. Szentaskó, Veresegyház, Hungary); 12.74, 10.7 (T. Vanmunster, Landen, Belgium).

1995 January 13 Brian G. Marsden